OpenStudy (anonymous):

-3x+2y=-4 Find the value of y when x=0 ;-;

OpenStudy (anonymous):

2y=-4

OpenStudy (anonymous):

y=-2

OpenStudy (catlover5925):



OpenStudy (anonymous):

-3x + 2y = -4 Just plug in x. -3(0) + 2y = -4 0 + 2y = -4 2y = -4 Divide each side by 2. y = -2 <---- Answer. (: